### v2

# SEFL: Enhancing Educational Assignment Feedback with LLM Agents

Abstract
Providing high-quality feedback to student assignments is crucial for student success, but it is constrained by time and costs. In this work, we introduce Synthetic Educational Feedback Loops (SEFL), a synthetic data framework designed to generate data that resembles immediate, on-demand feedback at scale without relying on extensive, real-world student assignments. To get this type of data, two large language models (LLMs) operate in teacher–student roles to simulate assignment completion and formative feedback, generating synthetic pairs of student work and corresponding critiques and actionable improvements from a teacher. With this data, we fine-tune smaller, more computationally efficient LLMs on these synthetic pairs, enabling them to replicate key features of high-quality, goal-oriented feedback. Unlike personalized tutoring approaches that offer multi-turn, individualized instruction, SEFL specifically focuses on replicating the teacherstudent assignment feedback loop in higher education. Through comprehensive evaluations with four LLM judges and three human experts, we demonstrate that SEFL-tuned models outperform both their non-tuned counterparts in feedback quality and an existing baseline. The potential for societal impact is reinforced by extensive qualitative comments by ratings by human stakeholders—both students and higher education instructors. All in all, SEFL has substantial potential to transform feedback processes for higher education and beyond.
Code — https://github.com/jjzha/sefl
Datasets and Models — https://huggingface.co/collections/jjzha/sefl-synthetic-educational-feedback-loops-67b48768dab5123a2e3e9d69
1 Introduction
Constructive feedback is a cornerstone of higher education, promoting critical thinking and fostering deeper understanding (Hattie 2008; Costello and Crane 2013). In higher education settings, however, providing consistent, high-quality feedback is complicated by privacy, consent, and transparency considerations in data collection (Fischer et al. 2020; Suresh et al. 2022; Demszky and Hill 2023; Wang and Demszky 2024; Wang et al. 2024b; Lindsay et al. 2024), in addition to being a labor-intensive task. Advances in language technology offer opportunities to automate and augment feedback processes, addressing these limitations.
In particular, LLMs have shown progress in education (Wang et al. 2024c), including automated grading (Ke and Ng 2019; Ramesh and Sanampudi 2022; Stahl et al. 2024) and personalized tutoring (Yun et al. 2024; Liu et al. 2024c; Rooein and Hovy 2024; Ross and Andreas 2024; Kwon et al. 2024; Zhang et al. 2024, 2025; Wang et al. 2024a). Yet, automating teacher–student assignment feedback with LLMs remains an open question. We seek to answer: RQ. How can synthetic teacher–student interactions generated by LLMs be leveraged to enable scalable and effective educational student assessment feedback?
Here, we introduce Synthetic Educational Feedback Loops (SEFL), a framework that generates synthetic teacher–student interactions using LLM agents. In this framework, two LLMs, one acting as the teacher and the other as the student, simulating formative feedback workflows (Conole and Oliver 2006; Nicol 2007). To induce better feedback from LLMs this synthetic data from the agents is then used to fine-tune smaller autoregressive language models, allowing the development of scalable educational feedback systems that can operate efficiently on modest computational infrastructure, available in higher education institutions without the need for privacy sensitive data.
We present several findings. (i) Through empirical and qualitative analysis of expert annotator comments, we find that larger models tend to provide more actionable, goal-oriented, and user-friendly feedback. They are also more consistent and better at supporting student autonomy, showing the potential for use in real-world contexts. (ii) Empirically, SEFL-tuned models outperform their non-tuned versions in win rate evaluations by 4 LLM judges and 3 human experts in giving assessment feedback. (iii) We observe strong agreement among human evaluators regarding feedback quality. (iv) We compare our approach to Book2Dial (Wang et al. 2024a), showing that SEFL achieves a higher win rate according to 3 of the 4 LLM judges.
Contributions.
We contribute the following in this work:
SEFL: A framework that simulates teacher student feedback loops with paired language model agents.
19,841 assignment–feedback pairs generated by SEFL to fine-tune smaller language models.
A comprehensive mix of human and language model evaluation that highlights strengths and limitations of SEFL with extensive qualitative analysis of the feedback provided by human experts.
An open source release of all models, code, and data.
2 Related Work
NLP & Education.
Large language models are now supporting a broad spectrum of educational tasks. In automated grading, they score short answers, essays, and even programming assignments with accuracy that approaches expert instructors, which eases faculty workload and releases time for mentoring (Ke and Ng 2019; Ramesh and Sanampudi 2022; Stahl et al. 2024). For personalized tutoring, conversational agents powered by these models adapt explanations, hints, and examples to each learner’s background knowledge and preferred style, producing measurable gains in engagement and achievement (Yun et al. 2024; Liu et al. 2024c; Rooein and Hovy 2024; Ross and Andreas 2024; Kwon et al. 2024; Zhang et al. 2024). Research on peer learning shows that the same technology can mediate small-group discussions, suggest prompts, and highlight diverse viewpoints, leading to richer collaboration (Bauer et al. 2023). In mathematics, aligning word problems and proofs to grade-level objectives has been automated with encouraging results, reducing the manual effort needed to curate question banks (Botelho et al. 2023). Critical thinking curricula benefit as well; LLMs can challenge students to justify claims, detect fallacies, and refine arguments in real time (Guerraoui et al. 2023). The models have also begun to assist scholars: studies report successful use for screening literature, summarizing drafts, and aligning reviewer comments with revision plans (Liang et al. 2024; Sonkar et al. 2024). Complementing these functions are analytics tools that track learning trajectories and surface early warnings when a student slips behind (Schwarz et al. 2018; Aslan et al. 2019; Alrajhi et al. 2021).
Despite this growing body of work, prior studies have rarely targeted the systematic generation of rich feedback on open-ended student submissions and assignments at scale. The present study addresses that gap by using LLMs to create extensive comment sets that teachers can accept as-is or trim to taste. Decades of scholarship define effective feedback as goal-oriented, actionable, timely, user-friendly, and consistent while fostering self-evaluation (Carless et al. 2011; Wiggins 2012). Feedback that rambles can confuse learners, so concise wording is usually preferable, and comments that arrive soon after the original effort drive steady improvement (Wiggins 2012). By producing immediate responses and tailoring suggestions to rubric criteria, LLM based systems stand well positioned to satisfy these guidelines while operating at classroom and institution scale.
Synthetic Data Frameworks.
Recent research shows how collaborative agentic LLMs can synthesize large-scale interactional datasets for educational tasks. For example, CAMEL (Li et al. 2023) uses cooperative role-based dialogues to achieve shared objectives, while SimSeek (Kim et al. 2022) uses agent-based conversations to build comprehensive information-seeking datasets. In education, SocraticLM (Liu et al. 2024b) simulates Socratic tutoring through multi-turn dialogue, and Book2Dial (Wang et al. 2024a) generates teacher-student conversations from textbooks. In contrast, SEFL focuses on concise teacher-student feedback loops rather than extended instructional dialogues. While Nair et al. (2024) explore iterative revisions, SEFL generates diverse feedback pairs from assignment-answer-feedback tuples, enabling fine-tuning of smaller, cost-effective models for large-scale use.
3 Synthetic Educational Feedback Loops
3.1 Synthetic Data Generation
We use a two-agent framework (Wu et al. 2023) to simulate a feedback loop as in higher education. Both the teacher and student roles are simulated by two separate Llama-3.1-70B models for a two-turn conversation.111Note that if we mention a model, it is always the post-trained version (i.e., -Instruct). The models are tasked to generate assignmentanswerfeedback tuples. First, the student-agent asks for an assignment using Fineweb-Edu (Lozhkov et al. 2024) texts (Figure 1), known for its educational content based on LLM judgments. Second, the teacher-agent creates an assignment that can be of any domain, e.g., math, humanities, role-playing. Then, the student-agent ( ) submits assignments containing a number of explicit errors, and the teacher-agent ( ) provides feedback. We investigate both Qwen2.5-72B and Llama-3.1-70B for interactions. First, we generate 5,000 interaction tuples with each model, where we validate the output as an initial step to investigate initial feedback quality.
We show in Table 1 the results of this first experiment. Out of 5,000 generated examples, Llama-3.1-70B generates 2,513 valid examples (i.e., valid JSON format and each feedback point refers to an error) compared to Qwen2.5-72B with 454 valid examples. For a further check, we use BERTScore (Zhang et al. 2020) as a proxy to see whether each error-feedback pair of the valid generations relate to each other.222We only calculate it of the samples where both error and feedback have the same number of generations. We show that, regardless of Llama-3.1-70B generating more valid examples, the BERTScore (0.877) stays in a similar range as Qwen2.5-72B (0.919); in both cases indicating high similarity. As a last qualitative check, we experimented with several prompts and consolidated into a final prompt found in the supplementary material. Finally, we use Llama-3.1-70B-generated data as the basis for all subsequent model fine-tuning as it generated more valid examples.
| Valid (/ 5,000) | BERTScore | |
| Llama-3.1-70B | 2,513 | 0.877 |
| Qwen2.5-72B | 454 | 0.919 |
Statistics.
After the 5K generated examples, we continue generating example pairs and end up with 19.8K teacher–student feedback pairs. In Table 2, we present the final dataset statistics. We highlight that the generation lengths for each agent are intentionally kept concise (170 subword tokens), based on the hypothesis that overly lengthy feedback may be counterproductive, especially for short assignments. This is in line with observations from Ferguson (2011), who finds that students tend to favor brief comments. We argue that balancing supportive and critical feedback is crucial as, by default, LLMs often produce excessively verbose responses, which can influence the preferences of both humans and language models (Saito et al. 2023).
4 Methodology
4.1 Fine-Tuning Large Language Models
We divide the data into 17,856 training examples and 1,985 validation examples. To test the generalizability of our approach across model scales, we fine-tune five open-weight models, namely Qwen2.5-0.5B, Llama-3.2-1B, Llama-3.2-3B, Llama-3.1-8B, and Qwen2.5-14B on this corpus. The compute we train the models on are AMD Radeon Instinct MI250X GPUs and it took a total of 467 GPU hours.
Supervised Objective.
To fine-tune the models, formally, for each prompt and target sequence we minimize the token-level cross-entropy
where masks out the prompt tokens and activates the loss on reference tokens only.
Optimization Details.
All models train for three epochs with a global batch size of 16 and context lengths of 131K for Qwen2.5 and 128K for the Llama variants. We use AdamW with , (Loshchilov and Hutter 2019), , weight decay , and gradient clipping at norm . The learning rate peaks at after a linear warm-up covering the first five percent of steps and then follows a linear decay.
| Feature | Value |
| Instances | 19,841 |
| Assignment Length | 78.6 |
| Length (Student Agent) | 168.1 |
| # Errors Points | 2.5 |
| Length # Errors | 20.7 |
| Length (Teacher Agent) | 120.5 |
| # Feedback Points | 2.5 |
| Length # Feedback | 34.6 |
4.2 Multi-faceted Evaluation
Human Evaluation.
After fine-tuning the language models with the assignment–feedback pairs, to test the performance of SEFL, we do human evaluation. We randomly sample 150 instances from the validation set. For each item, both the original instruction tuned model (A) and the model further fine-tuned with SEFL (B) produce feedback. Three human experts compared pairs of feedback responses produced for the same assignment and answer. For each item they read the original prompt, the student submission, and the two candidate feedback drafts from the non-tuned and SEFL-tuned model. Then, they select the feedback, from model A or B, that is better based on four base criteria:
Accuracy. The generated feedback text refers to concrete strengths and weaknesses in the student work and avoid superficial remarks.
Actionability. Suggestions are clear, specific, and realistic for a student to apply.
Conciseness. Wording is brief and focused, with little repetition.
Tone. Language stays constructive and professional while recognizing good elements.
Raters were reminded to value efficiency over length, to prefer targeted advice over general principles, and to ignore formatting tricks. They recorded their choice as A or B and could leave an optional free-text comment. With this we calculate win rate (i.e., the percentage of choosing one feedback text over the other). This has become a de facto standard to evaluate long form text against each other (e.g., Rafailov et al. 2023).
Each row took at most ten minutes, and the guidelines stressed taking regular breaks to sustain attention. We deliberately remove the tie option because a forced choice gives more informative labels and reduces hesitation, while a separate checkbox still lets raters mark assignmentanswerfeedback tuples as unrelated or nonsensical. This happened around 12% of the time, especially in the smaller, less capable models. The full annotation guidelines reported in the supplementary material in the code repository.
Our human raters are in the aged 20–40 and from Europe. One identifies as female and the other two identifies as male. One female and male have a background in Computer Science and one male in Education Engineering. All have extensive experience in teaching and supervision or being taught and supervised, they all work in higher education (at different levels, i.e., research assistant and assistant professors) with near-native English proficiency.
LLM-as-a-Judge.
We also evaluate the fine-tuned models’ output using a LLM-as-a-judge framework, a method gaining traction as a method for evaluating free text output (Liu et al. 2023; Zheng et al. 2024; Chen et al. 2023; Verga et al. 2024; Törnberg 2023; Naismith, Mulcaire, and Burstein 2023; Gilardi, Alizadeh, and Kubli 2023; Kocmi and Federmann 2023; Huang et al. 2024; Gu et al. 2024; Falk et al. 2025). The same 150 random instances are rated by the four LLMs, namely GPT-4o (Hurst et al. 2024), Claude3.5-Sonnet, Command-R+, and DeepSeek-V3 (Liu et al. 2024a). For the closed-source models’ LLM-as-a-judge experiments, we use their respective APIs and the total costs were approximately 10 USD.
For every example the judge model receives the assignment prompt together with the two candidate feedback drafts. It is asked to decide which draft is better on the same four base criteria as the humans. The instruction forbids numeric grades or explanations and requires the judge to output exactly one character, A or B, producing a clean pairwise preference label. We report the full prompt in the supplementary material in the code repository.
5 Results
In Figure 2, we show the win rates of models fine-tuned with SEFL vs. their non-tuned version, evaluated by both humans and LLM-based judges. A value above 50% indicates that the SEFL-tuned models are preferred over their original versions. We show an example of the feedback in Figure 4, where we depict the abridged prompt and feedback by two tuned an non-tuned models.
Human Assessment.
Overall, human rater evaluations in Figure 2 (H#) show that the SEFL-tuned models often attain high win rates, surpassing 90% in several cases with respect to the smaller models. The human annotators differed in their views on the 8B model’s output quality; however, they generally converged on the observation that the fine-tuned 14B model produces superior feedback compared to its original version. By contrast, models not fine-tuned with SEFL had lower win rates, suggesting that SEFL provides an edge in generating more coherent and context-relevant feedback. In addition, we asked annotators whether the synthetic assignmentanswerfeedback sequences were consistent. In over 75% of cases, they confirmed the alignment between assignment, student response, and the feedback given, showing positive contextual relevance.
LLM-as-a-Judge Results.
For the LLM-as-a-judge evaluations (J#) in the same figure, we observe notable differences in win rates depending on the model and scale. The results largely mirror the human assessment trend up to the 3B scale. The results from the four LLM judges (J1: GPT-4o, J2: Claude-3.5-Sonnet, J3: Command-R+, J4: Deepseek-v3) reveal that SEFL-tuned models show varying levels of performance relative to their vanilla counterparts.333Models are picked based on their recency and performance on RewardBench (Lambert et al. 2024), JudgeBench (Tan et al. 2024), and JudgeArena (AtlaAI 2025). For instance, Qwen2.5-0.5B achieved the highest win rates across all four judges (62% on J3), indicating a consistent preference for the fine-tuned version. In contrast, larger models such as Llama-3.1-8B and Qwen2.5-14B exhibit lower win rates, particularly on J3 (16% and 10%, respectively), suggesting that fine-tuning with SEFL may yield diminishing returns or challenges at larger scales. The full judge prompt can be found in the supplementary material.
Agreement.
In Figure 3, we present the pairwise Cohen’s values (Cohen 1960) computed between each language-model judge and the human raters, to observe whether humans and LLM judges agree on which model gives better feedback. The agreement among humans was moderate to substantial: H1 and H3 reached , H1 and H2 , and H2 and H3 (Landis and Koch 1977). Among the models, Claude aligns most closely with both the other judges and the humans; DeepSeek follows, and GPT-4o shows the weakest match. Across all model and human pairs the coefficients range from to , which underlines the subjectivity of feedback evaluation. Command R+ sits at the lower extreme, returning values between and when compared with human raters and with other judges, which signals virtually no agreement. This signals that the human experts mainly agree on the quality of feedback, but for LLM judges this can differ. Overall, there are improvements to be made in terms of agreement between LLMs and humans in terms of feedback quality.
[…]
| 0.5B | “The feedback repeats the assignment rather than pointing out issues.” | “A is concise but misses important detail.” | “Comments do not reflect missing parts of the answer.” |
| 0.5B-SEFL | “Feedback spots deliberate errors and offers clear fixes.” | “A is concise and clear, suggestions align with the rubric.” | “Advice notes that the discuss part is missing and proposes adding it.” |
| 1B | “Feedback echoes what the answer already states.” | “More precise.” | “Comments ignore omitted sections.” |
| 1B-SEFL | “The note opens with strengths and gives concrete next steps.” | “Hints link directly to rubric points.” | “Advice adds two practical examples to guide revision.” |
| 3B | “Both models are good, A sounds nicer.” | “Both are good here.” | “Feedback covers content but wording feels harsh.” |
| 3B-SEFL | “Feedback keeps an encouraging tone while giving actionable points.” | “Structure and tone feel balanced.” | “Comments cover content, style, and structure in one cohesive note.” |
| 7B | “Feedback addresses surface errors but misses deeper reasoning.” | “Detailed but drifts off the prompt.” | “Overlooks the rubric item about evidence.” |
| 7B-SEFL | “Points out reasoning gaps and proposes specific fixes.” | “Stays focused on the prompt and remains concise.” | “Directly cites the missing evidence section.” |
| 14B | “Feedback is thorough but wording is opaque.” | “Long sentences reduce clarity.” | “Some comments repeat earlier points.” |
| 14B-SEFL | “Clear and student friendly while retaining depth.” | “Uses short sentences and connects advice to rubric.” | “Groups comments by theme which avoids repetition.” |
| Aspect | Base Model | SEFL-tuned Model | Preferred |
| Consistency | “Feedback from model A fails to address key aspects of the answer, such as suddenly changing the name of the main character.” | “Not an answer but A properly identified it!” | SEFL |
| Goal orientation | “In many cases, answers are shorter than required. Not reflected in feedback.” | “Finally, B finds that the answer is incomplete.” | SEFL |
| Tone | “Model B: Harsh tone.” | “Both models are good, but model A is nicer in tone and actionability.” | SEFL |
| User friendliness | “Model B is best, but is way too elaborate.” | “B is more clear and concise.” | SEFL |
| Actionability | “Model A also provides partial solution.” | “Model B is not accurate and provides an answer instead of feedback.” | Base |
6 Discussion
6.1 Human Qualitative Insights
In addition to the win rates in Figure 2, our human annotators provided rich qualitative feedback on the model outputs, which we show in Table 3. Generally, on the critical side, they noted that if a student answer is too short or incomplete, neither model explicitly flags the missing details. More specifically, Qwen2.5-0.5B was praised for clarity and concision, whereas Llama-3.2-3B tended to repeat assignment details without offering actionable guidance. Annotators observed that Llama-3.2-1B often gave more specific and constructive feedback but occasionally sounded harsh, while Llama-3.1-8B sometimes overlooked key aspects. Overall, although Qwen2.5-14B achieved high win rates (94, 77, 81 across three annotators), these insights suggest that even top-performing models could improve in error detection, tone refinement, and contextual sensitivity.
To further quantify, in Figure 5, we plot for each model the net balance of optional rater comments in five qualitative categories: Actionability (AC), Goal Orientation (GO), User Friendliness (UF), Consistency (CO) and Student Autonomy (AY) (Carless et al. 2011; Wiggins 2012). Squares denote the base models and circles the SEFL-tuned variants. Horizontal whiskers give bootstrap Wilson 95% confidence intervals for the net balance. We compute these intervals on the proportion of positive remarks and then transform them to the net scale via , where is the total number of comments and the positive proportion. Annotators added comments only when a response stood out, so the plot reveals both the direction and the strength of impressions. We show several examples of annotations in Table 4. The annotators also had a choice if the feedback was non-nonsensical, which happened around 12% of the time mainly with the smaller sized models.
The two smallest models, Qwen2.5-0.5B and Llama-3.2-1B, receive more negative than positive remarks on Consistency and Goal Orientation, matching earlier findings that they sometimes drift from the student answer or overlook core requirements. Llama-3.1-8B shows the only clearly positive balance in Actionability and User Friendliness, yet its interval for Consistency still lies below zero. Qwen2.5-14B gains more favorable notes on tone and clarity than the smaller models but still shows a negative alignment gap. With respect to Student Autonomy, there are mostly neutral comments.
| Method/Dataset | J1 | J2 | J3 | J4 |
| Book2Dial | 43 | 37 | 60 | 27 |
| SEFL | 57 | 63 | 40 | 73 |
6.2 LLM-as-a-Judge
We used LLM judges to rate the feedback generated by SEFL-tuned models against their vanilla counterparts. This provides a rapid, scalable way to measure feedback quality, reducing the need for extensive human annotation. We let the LLM judges rate the same examples as the humans annotated. As shown in Figure 2, three out of four LLM judges consistently favored SEFL-tuned Qwen2.5-0.5B, Llama-3.2-1B, and Llama-3.2-3B. With Command-R, we notice that it performs worse than GPT-4o and Claude-3.5-Sonnet on JudgeArena. This indicates that the misalignment might have to do with instruction following capabilities of these models, this is highlighted by Command-R being weaker in Chatbot Arena (Zheng et al. 2024). Nonetheless, we see it as a practical first step for large-scale feedback comparisons in educational contexts. We recommend supplementing LLM-based assessments with targeted human evaluations for more granular insights, possibly aligning more with instructional objectives or possibly even fine-tuning LLMs with rubrics for better judges of long form text (Kim et al. 2024).
6.3 Comparison to Prior Work
The work that is closest to ours is Book2Dial (Wang et al. 2024a). It is a framework that turns textbooks into synthetic conversations between a student model and a teacher model. The student only sees high-level cues such as section titles or key concepts, while the teacher has full access to the source passage, prompting a question-answer exchange that stays aligned with the book content. Instead, we focus on any type of assignment not limited to textbooks.
To compare the two methods, we further fine-tune Qwen2.5-14B with the Book2Dial data in the same way as our SEFL-tuned version. We use the existing Book2Dial data consisting of 889 dialogues, we preprocess the data such that there is only a one-turn conversation, resulting in 5.3K conversation pairs. We then run the fine-tuned model over the same samples evaluated in Subsection 4.2 and evaluate them with the same 4 judges and evaluation criteria.444At the time of writing, we ran out of funding for the human annotators and thus compare here only with LLM-as-a-Judge.
Table 5 shows that three out of four judges prefer SEFL, yielding an average win rate of 58%. Only command-r-plus leans the other way with a 60% win rate for Book2Dial, but an average of 42% win rate. The finding is consistent with its earlier drift of seemingly failing to instruction follow. The results confirm that SEFL produces higher quality feedback than the textbook based dialogues in Book2Dial.
7 Conclusion
We introduced SEFL, a framework that simulates teacherstudent interactions via two-agent LLMs to generate synthetic data for fine-tuning smaller models. Empirical and qualitative results show the potential of this framework for enhancing feedback in educational settings. This approach yields concise, context-sensitive feedback that often surpasses original instruction-tuned models under both LLM-as-a-judge and human evaluations. Yet human insights remain indispensable for capturing nuances like clarity and tone. SEFL provides a promising avenue for immediate, personalized feedback at scale, extending beyond the educational domain.
Limitations.
We acknowledge that SEFL relies on synthetically generated assignments and errors, and are not real student submissions, which could have implications. Although this approach helps create large datasets, it risks producing feedback unaligned with authentic classroom contexts. Our evaluation also uses LLM-based judges, introducing potential biases related to each judge’s training data and objectives. Lastly, while we focused on short-answer tasks, longer or more domain-specific assignments may require specialized or more diverse synthetic data.
Ethical Statement
The use of synthetic data provides an opportunity to train automated feedback systems without the constraints of privacy and consent that come from repurposing actual student assignments and teacher feedback as training data. However, it also raises questions about transparency and potential misuse (Lindsay et al. 2024). For instance, malicious actors could manipulate synthetic data to disseminate misleading or biased feedback, undermining trust in educational tools. Users may also mistake synthetic feedback for real, expert guidance. Moreover, automated feedback systems risk reinforcing biases if the underlying models carry skewed training data. We believe educators and institutions should remain aware of these risks and incorporate human oversight to ensure that such systems complement, rather than replace, genuine pedagogical engagement with real teachers.

### v3

@@
-SEFL: Enhancing Educational Assignment Feedback with LLM Agents
+SEFL: A Framework for Generating Synthetic Educational Assignment Feedback with LLM Agents
 Abstract
-Providing high-quality feedback to student assignments is crucial for student success, but it is constrained by time and costs.
-In this work, we introduce Synthetic Educational Feedback Loops (SEFL), a synthetic data framework designed to generate data that resembles immediate, on-demand feedback at scale without relying on extensive, real-world student assignments.
-To get this type of data, two large language models (LLMs) operate in teacher–student roles to simulate assignment completion and formative feedback, generating synthetic pairs of student work and corresponding critiques and actionable improvements from a teacher.
+Providing high-quality feedback on student assignments is crucial for student success, but it is heavily limited by time and budgetary constraints.
+In this work, we introduce Synthetic Educational Feedback Loops (SEFL), a synthetic data framework designed to generate data that resembles immediate, on-demand feedback at scale without relying on extensive, real-world student assignments and teacher feedback.
+To obtain this type of data, two large language models (LLMs) operate in a teacher-student role to simulate assignment completion and formative feedback, generating 19.8K synthetic pairs of student work and corresponding critiques and actionable improvements from a teacher.
 With this data, we fine-tune smaller, more computationally efficient LLMs on these synthetic pairs, enabling them to replicate key features of high-quality, goal-oriented feedback.
-Unlike personalized tutoring approaches that offer multi-turn, individualized instruction, SEFL specifically focuses on replicating the teacherstudent assignment feedback loop in higher education.
-Through comprehensive evaluations with four LLM judges and three human experts, we demonstrate that SEFL-tuned models outperform both their non-tuned counterparts in feedback quality and an existing baseline.
-The potential for societal impact is reinforced by extensive qualitative comments by ratings by human stakeholders—both students and higher education instructors.
-All in all, SEFL has substantial potential to transform feedback processes for higher education and beyond.
-Code — https://github.com/jjzha/sefl
-Datasets and Models — https://huggingface.co/collections/jjzha/sefl-synthetic-educational-feedback-loops-67b48768dab5123a2e3e9d69
-1 Introduction
-Constructive feedback is a cornerstone of higher education, promoting critical thinking and fostering deeper understanding (Hattie 2008; Costello and Crane 2013).
-In higher education settings, however, providing consistent, high-quality feedback is complicated by privacy, consent, and transparency considerations in data collection (Fischer et al.
-2020; Suresh et al.
-2022; Demszky and Hill 2023; Wang and Demszky 2024; Wang et al.
-2024b; Lindsay et al.
-2024), in addition to being a labor-intensive task.
-Advances in language technology offer opportunities to automate and augment feedback processes, addressing these limitations.
-In particular, LLMs have shown progress in education (Wang et al.
-2024c), including automated grading (Ke and Ng 2019; Ramesh and Sanampudi 2022; Stahl et al.
-2024) and personalized tutoring (Yun et al.
-2024; Liu et al.
-2024c; Rooein and Hovy 2024; Ross and Andreas 2024; Kwon et al.
-2024; Zhang et al.
-2024, 2025; Wang et al.
-2024a).
-Yet, automating teacher–student assignment feedback with LLMs remains an open question.
+Through comprehensive evaluations with three LLM judges and three human experts, across a subset of 900 outputs, we demonstrate that SEFL-tuned models outperform both their untuned counterparts and an existing baseline in terms of feedback quality.
+The potential for societal impact is reinforced by extensive qualitative comments and ratings from human stakeholders — both students and higher education instructors.
+SEFL has the potential to transform feedback processes for higher education and beyond.
+Keywords: Synthetic Data, Education, Large Language Model, Agents
+SEFL: A Framework for Generating Synthetic Educational Assignment Feedback with LLM Agents
+| Mike Zhang, Amalie Pernille Dilling, Léon Gondelman |
+| Niels Erik Ruan Lyngdorf, Euan D. Lindsay, Johannes Bjerva |
+| University of Copenhagen, Denmark |
+| Aalborg University, Denmark |
+| Pioneer Centre for AI, Denmark |
+| mike.zhang@di.ku.dk |
+Abstract content
+1. Introduction
+Constructive feedback is a cornerstone of higher education, promoting critical thinking and fostering deeper understanding Hattie (2008); Costello and Crane (2013).
+In higher education settings, however, providing consistent, high-quality feedback is complicated by privacy, consent, and transparency considerations in data collection Fischer et al.
+(2020); Suresh et al.
+(2022); Demszky and Hill (2023); Wang and Demszky (2024); Wang et al.
+(2024b); Lindsay et al.
+(2024), in addition to being a labor-intensive task.
+Advances in language technology offer opportunities to automate and augment higher education feedback processes, addressing these limitations.
+In particular, LLMs have shown progress in education Wang et al.
+(2024c), including automated grading Ke and Ng (2019); Ramesh and Sanampudi (2022); Stahl et al.
+(2024) and personalized tutoring Yun et al.
+(2024); Liu et al.
+(2024c); Rooein and Hovy (2024); Ross and Andreas (2024); Kwon et al.
+(2024); Zhang et al.
+(2024, 2025); Wang et al.
+(2024a).
+Yet, automating teacher-student assignment feedback with LLMs remains an open question.
 We seek to answer: RQ.
-How can synthetic teacher–student interactions generated by LLMs be leveraged to enable scalable and effective educational student assessment feedback?
-Here, we introduce Synthetic Educational Feedback Loops (SEFL), a framework that generates synthetic teacher–student interactions using LLM agents.
-In this framework, two LLMs, one acting as the teacher and the other as the student, simulating formative feedback workflows (Conole and Oliver 2006; Nicol 2007).
-To induce better feedback from LLMs this synthetic data from the agents is then used to fine-tune smaller autoregressive language models, allowing the development of scalable educational feedback systems that can operate efficiently on modest computational infrastructure, available in higher education institutions without the need for privacy sensitive data.
-We present several findings.
-(i) Through empirical and qualitative analysis of expert annotator comments, we find that larger models tend to provide more actionable, goal-oriented, and user-friendly feedback.
-They are also more consistent and better at supporting student autonomy, showing the potential for use in real-world contexts.
-(ii) Empirically, SEFL-tuned models outperform their non-tuned versions in win rate evaluations by 4 LLM judges and 3 human experts in giving assessment feedback.
+How can synthetic teacher-student interactions generated by LLMs be leveraged to enable scalable and effective educational student assessment feedback?
+Here, we introduce Synthetic Educational Feedback Loops (SEFL), a framework that generates synthetic teacher-student interactions using LLM agents.
+In this framework, two LLMs, one acting as the teacher and the other as the student, simulate formative feedback workflows Conole and Oliver (2006); Nicol (2007).
+This synthetic data from the agents is then used to fine-tune smaller autoregressive language models, resulting in better feedback models that enable the development of scalable educational feedback systems.
+These systems can operate efficiently on the modest computational infrastructure available in higher education institutions, without requiring access to privacy-sensitive data.
+We show that: (i) Empirically, SEFL-tuned models outperform their non-tuned versions in win rate evaluations by three LLM judges and three human experts in giving assessment feedback.
+(ii) Through empirical and qualitative analysis of expert annotator comments, we find that larger models tend to provide more actionable, goal-oriented, and user-friendly feedback.
+They are also more consistent and better at supporting student autonomy.
 (iii) We observe strong agreement among human evaluators regarding feedback quality.
-(iv) We compare our approach to Book2Dial (Wang et al.
-2024a), showing that SEFL achieves a higher win rate according to 3 of the 4 LLM judges.
+(iv) We compare our approach to an existing baseline (Book2Dial; Wang et al., 2024a) and show that SEFL-tuned models provide better feedback according to all LLM judges.
 Contributions.
-We contribute the following in this work:
-SEFL: A framework that simulates teacher student feedback loops with paired language model agents.
-19,841 assignment–feedback pairs generated by SEFL to fine-tune smaller language models.
-A comprehensive mix of human and language model evaluation that highlights strengths and limitations of SEFL with extensive qualitative analysis of the feedback provided by human experts.
-An open source release of all models, code, and data.
-2 Related Work
+We contribute the following:
+SEFL: An agentic framework that simulates teacher-student feedback loops with paired language model agents.
+19,841 synthetic assignment–feedback pairs generated by SEFL to fine-tune smaller language models.
+A comprehensive set of human and LLM evaluations that outline the strengths and limitations of SEFL with extensive qualitative analysis of the feedback provided by human experts.
+Open-sourcing of all models, code, and data.
+111Resources and Supplementary Material such as full prompts and example outputs can be found at https://github.com/jjzha/sefl and https://tinyurl.com/3zdu847k (HuggingFace).
+2. Related Work
 NLP & Education.
 Large language models are now supporting a broad spectrum of educational tasks.
-In automated grading, they score short answers, essays, and even programming assignments with accuracy that approaches expert instructors, which eases faculty workload and releases time for mentoring (Ke and Ng 2019; Ramesh and Sanampudi 2022; Stahl et al.
-2024).
-For personalized tutoring, conversational agents powered by these models adapt explanations, hints, and examples to each learner’s background knowledge and preferred style, producing measurable gains in engagement and achievement (Yun et al.
-2024; Liu et al.
-2024c; Rooein and Hovy 2024; Ross and Andreas 2024; Kwon et al.
-2024; Zhang et al.
-2024).
-Research on peer learning shows that the same technology can mediate small-group discussions, suggest prompts, and highlight diverse viewpoints, leading to richer collaboration (Bauer et al.
-2023).
-In mathematics, aligning word problems and proofs to grade-level objectives has been automated with encouraging results, reducing the manual effort needed to curate question banks (Botelho et al.
-2023).
-Critical thinking curricula benefit as well; LLMs can challenge students to justify claims, detect fallacies, and refine arguments in real time (Guerraoui et al.
-2023).
-The models have also begun to assist scholars: studies report successful use for screening literature, summarizing drafts, and aligning reviewer comments with revision plans (Liang et al.
-2024; Sonkar et al.
-2024).
-Complementing these functions are analytics tools that track learning trajectories and surface early warnings when a student slips behind (Schwarz et al.
-2018; Aslan et al.
-2019; Alrajhi et al.
-2021).
-Despite this growing body of work, prior studies have rarely targeted the systematic generation of rich feedback on open-ended student submissions and assignments at scale.
-The present study addresses that gap by using LLMs to create extensive comment sets that teachers can accept as-is or trim to taste.
-Decades of scholarship define effective feedback as goal-oriented, actionable, timely, user-friendly, and consistent while fostering self-evaluation (Carless et al.
-2011; Wiggins 2012).
-Feedback that rambles can confuse learners, so concise wording is usually preferable, and comments that arrive soon after the original effort drive steady improvement (Wiggins 2012).
-By producing immediate responses and tailoring suggestions to rubric criteria, LLM based systems stand well positioned to satisfy these guidelines while operating at classroom and institution scale.
+In automated grading, they score short answers, essays, and even programming assignments with accuracy that approaches expert instructors, which eases faculty workload and releases time for mentoring Ke and Ng (2019); Ramesh and Sanampudi (2022); Stahl et al.
+(2024).
+For personalized tutoring, conversational agents powered by these models adapt explanations, hints, and examples to each learner’s background knowledge and preferred style, producing measurable gains in engagement and achievement Yun et al.
+(2024); Liu et al.
+(2024c); Rooein and Hovy (2024); Ross and Andreas (2024); Kwon et al.
+(2024); Zhang et al.
+(2024).
+Research on peer learning shows that the same technology can mediate small-group discussions, suggest prompts, and highlight diverse viewpoints, leading to richer collaboration Bauer et al.
+(2023).
+In mathematics, aligning word problems and proofs to grade-level objectives has been automated, yielding encouraging results and reducing the manual effort required to curate question banks Botelho et al.
+(2023).
+Critical thinking curricula also benefit; LLMs can challenge students to justify claims, detect fallacies, and refine arguments in real-time Guerraoui et al.
+(2023).
+The models have also begun to assist scholars: studies report successful use for screening literature, summarizing drafts, and aligning reviewer comments with revision plans Liang et al.
+(2024); Sonkar et al.
+(2024).
+Complementing these functions are analytics tools that track learning trajectories and surface early warnings when a student slips behind Schwarz et al.
+(2018); Aslan et al.
+(2019); Alrajhi et al.
+(2021).
+Despite this growing body of work, prior studies have rarely targeted the systematic generation of feedback on open-ended student submissions and assignments.
+In this study, we address this gap by using LLMs to generate extensive comment sets that teachers can accept as is or adapt.
+Decades of scholarship define effective feedback as goal-oriented, actionable, timely, user-friendly, and consistent while fostering self-evaluation Carless et al.
+(2011); Wiggins (2012).
+Feedback that is long and verbose can confuse learners.
+Therefore, concise wording is usually preferable, and comments that arrive soon after the original effort drive steady improvement Wiggins (2012).
+By producing immediate responses and tailoring suggestions to rubric criteria, LLM-based systems stand well-positioned to satisfy these guidelines while operating at the classroom and institution scale.
 Synthetic Data Frameworks.
 Recent research shows how collaborative agentic LLMs can synthesize large-scale interactional datasets for educational tasks.
-For example, CAMEL (Li et al.
-2023) uses cooperative role-based dialogues to achieve shared objectives, while SimSeek (Kim et al.
-2022) uses agent-based conversations to build comprehensive information-seeking datasets.
-In education, SocraticLM (Liu et al.
-2024b) simulates Socratic tutoring through multi-turn dialogue, and Book2Dial (Wang et al.
-2024a) generates teacher-student conversations from textbooks.
+For example, CAMEL Li et al.
+(2023) employs cooperative role-based dialogues to achieve shared objectives, whereas SimSeek Kim et al.
+(2022) utilizes agent-based conversations to construct comprehensive information-seeking datasets.
+In education, SocraticLM Liu et al.
+(2024b) simulates Socratic tutoring through multi-turn dialogue, and Book2Dial Wang et al.
+(2024a) generates teacher-student conversations from textbooks.
 In contrast, SEFL focuses on concise teacher-student feedback loops rather than extended instructional dialogues.
@@
 (2024) explore iterative revisions, SEFL generates diverse feedback pairs from assignment-answer-feedback tuples, enabling fine-tuning of smaller, cost-effective models for large-scale use.
-3 Synthetic Educational Feedback Loops
-3.1 Synthetic Data Generation
-We use a two-agent framework (Wu et al.
-2023) to simulate a feedback loop as in higher education.
-Both the teacher and student roles are simulated by two separate Llama-3.1-70B models for a two-turn conversation.111Note that if we mention a model, it is always the post-trained version (i.e., -Instruct).
+3. SEFL: Synthetic Educational Feedback Loops
+3.1. Synthetic Data Generation
+We employ a two-agent framework Wu et al.
+(2023) to simulate a student-teacher feedback loop, as seen in higher education.
+Both the teacher and student roles are simulated by two separate Llama-3.1-70B models for a two-turn conversation.222Note that if we mention a model, it is always the post-trained version (i.e., -Instruct).
 The models are tasked to generate assignmentanswerfeedback tuples.
-First, the student-agent asks for an assignment using Fineweb-Edu (Lozhkov et al.
-2024) texts (Figure 1), known for its educational content based on LLM judgments.
-Second, the teacher-agent creates an assignment that can be of any domain, e.g., math, humanities, role-playing.
-Then, the student-agent ( ) submits assignments containing a number of explicit errors, and the teacher-agent ( ) provides feedback.
+First, the student-agent requests an assignment using Fineweb-Edu Lozhkov et al.
+(2024) texts (Figure˜1), which is known for its educational content based on LLM judgments.
+Second, the teacher-agent creates an assignment that can be of any domain, e.g., STEM, social sciences, puzzles, and so forth.
+Then, the student-agent ( ) submits assignments containing several explicit errors, and the teacher-agent ( ) provides feedback.
+In and , we obtain an assignment-feedback pair.
+We deliberately avoided imposing strict rules on the types of errors that could be generated.
+We hypothesize that allowing the LLM to introduce errors naturally results in a more diverse and realistic set of mistakes than if we had a fixed checklist.
 We investigate both Qwen2.5-72B and Llama-3.1-70B for interactions.
-First, we generate 5,000 interaction tuples with each model, where we validate the output as an initial step to investigate initial feedback quality.
-We show in Table 1 the results of this first experiment.
+For a control check, we generate 5,000 interaction tuples with each model, and validate the output as an initial step to investigate the quality of initial feedback.
+We show in Table˜1 the results of this experiment.
 Out of 5,000 generated examples, Llama-3.1-70B generates 2,513 valid examples (i.e., valid JSON format and each feedback point refers to an error) compared to Qwen2.5-72B with 454 valid examples.
-For a further check, we use BERTScore (Zhang et al.
-2020) as a proxy to see whether each error-feedback pair of the valid generations relate to each other.222We only calculate it of the samples where both error and feedback have the same number of generations.
-We show that, regardless of Llama-3.1-70B generating more valid examples, the BERTScore (0.877) stays in a similar range as Qwen2.5-72B (0.919); in both cases indicating high similarity.
-As a last qualitative check, we experimented with several prompts and consolidated into a final prompt found in the supplementary material.
-Finally, we use Llama-3.1-70B-generated data as the basis for all subsequent model fine-tuning as it generated more valid examples.
+For a further check, we use BERTScore Zhang et al.
+(2020) as a proxy to see whether each error-feedback pair of the valid generations relates to each other.333We only calculate it of the samples where both error and feedback have the same number of generations.
+We show that, regardless of Llama-3.1-70B generating more valid examples, the BERTScore (0.877) remains in a similar range to Qwen2.5-72B (0.919); in both cases, this indicates a high level of similarity.
+As a final qualitative check, we experimented with several prompts and consolidated them into a single, final prompt, which is included in the supplementary material.
+Finally, we use Llama-3.1-70B-generated data as the basis for all subsequent model fine-tuning, as it generated more valid examples.
+We spot-checked several prompts and consolidated the final full prompt in Figure˜6 (Appendix˜A).
 | Valid (/ 5,000) | BERTScore | |
@@
 | Qwen2.5-72B | 454 | 0.919 |
-Statistics.
-After the 5K generated examples, we continue generating example pairs and end up with 19.8K teacher–student feedback pairs.
-In Table 2, we present the final dataset statistics.
-We highlight that the generation lengths for each agent are intentionally kept concise (170 subword tokens), based on the hypothesis that overly lengthy feedback may be counterproductive, especially for short assignments.
+Data statistics.
+After generating 5,000 examples, we continue to create example pairs, resulting in 19,841 teacher-student feedback pairs.
+In Table˜3, we present the final dataset statistics.
+We underline that the generation lengths for each agent are intentionally kept concise (170 subword tokens), based on the hypothesis that overly lengthy feedback may be counterproductive.
 This is in line with observations from Ferguson (2011), who finds that students tend to favor brief comments.
-We argue that balancing supportive and critical feedback is crucial as, by default, LLMs often produce excessively verbose responses, which can influence the preferences of both humans and language models (Saito et al.
-2023).
-4 Methodology
-4.1 Fine-Tuning Large Language Models
+We argue that balancing supportive and critical feedback is crucial, as LLMs often produce verbose responses by default, which can influence the preferences of both humans and LMs Saito et al.
+(2023).
+Task Errors.
+Once all data was generated, we took a random subsample of 226 instances with 428 explicit errors.
+We investigated what category of intentional errors the student agent generates.
+In total, we identified seven error types from the 428 errors and show the proportion in Table˜2.
+Most errors are of type “Omission and/or incompleteness” (surmounted for 50% of the errors):
+1.
+Omission and/or incompleteness: This includes missing details, lacking examples or evidence, not exploring a topic deeply enough, or failing to address all parts of the assignment.
+2.
+Stylistic and/or formatting issues: Problems with grammar, spelling, punctuation, tone (e.g., too informal), word count, and incorrect formatting (like improper citations).
+3.
+Factual inaccuracy: This includes incorrect dates, numbers, names, scientific facts, historical events, or misquoting a source text.
+4.
+Logical flaws and/or weak argumentation: These are errors in reasoning.
+They include making incorrect assumptions, oversimplifying complex topics, drawing faulty conclusions, or failing to construct a coherent and well-supported argument.
+5.
+Technical and/or procedural errors: This category is for mistakes in technical execution, such as incorrect mathematical calculations, flawed experimental design, or errors in code.
+6.
+Structural and organizational problems: These errors relate to the overall structure and flow of the response, such as a lack of clear transitions between paragraphs, poor organization, or a missing introduction/conclusion.
+7.
+Task misinterpretation: Student agent misunderstands the core requirement of the task, such as writing a story instead of an essay or failing to use a requested framework.
+| Error Category | Count | % | Example Error Explanation from Student Agent |
+| Omission and/or Incompleteness | 214 | 50.0% | “Failure to provide specific examples and evidence to support claims”. |
+| Stylistic and/or Formatting Issues | 67 | 15.7% | “Incorrect spelling of the word ’protection’ in the last sentence of the answer”. |
+| Factual Inaccuracy | 59 | 13.8% | “Incorrectly claimed Ibn al-Haytham was born in Egypt instead of Basra, Iraq”. |
+| Logical Flaws and/or Weak Argumentation | 53 | 12.4% | “The statement about DSP chips having limitations is an oversimplification, as they are still widely used for various applications”. |
+| Technical and/or Procedural Errors | 16 | 3.7% | “Incorrect calculation of the target waiting time, should be a 30% reduction from 10 minutes”. |
+| Structural and/or Organizational Problems | 12 | 2.8% | “Lack of clear structure and organization in the comparative analysis”. |
+| Task Misinterpretation | 7 | 1.6% | “The writing deploys facts that were not mentioned in the letter, like United States’ interest in containing the spread of communism”. |
+| Total | 428 | 100.0% | — |
+4. Methodology
+4.1. Fine-Tuning Large Language Models
 We divide the data into 17,856 training examples and 1,985 validation examples.
-To test the generalizability of our approach across model scales, we fine-tune five open-weight models, namely Qwen2.5-0.5B, Llama-3.2-1B, Llama-3.2-3B, Llama-3.1-8B, and Qwen2.5-14B on this corpus.
-The compute we train the models on are AMD Radeon Instinct MI250X GPUs and it took a total of 467 GPU hours.
-Supervised Objective.
-To fine-tune the models, formally, for each prompt and target sequence we minimize the token-level cross-entropy
+To test the feedback quality of our approach across model scales, we fine-tune five open-weight models, namely Qwen2.5-0.5B, Llama-3.2-1B, Llama-3.2-3B, Llama-3.1-8B, and Qwen2.5-14B on this synthetic dataset.
+The compute we train the models on is AMD Radeon Instinct MI250X GPUs, and it took a total of 467 GPU hours.
+Training Objective.
+To fine-tune the LLMs, for each prompt and target sequence we minimize the token-level cross-entropy
 where masks out the prompt tokens and activates the loss on reference tokens only.
@@
 All models train for three epochs with a global batch size of 16 and context lengths of 131K for Qwen2.5 and 128K for the Llama variants.
-We use AdamW with , (Loshchilov and Hutter 2019), , weight decay , and gradient clipping at norm .
+We use AdamW with , Loshchilov and Hutter (2019), , weight decay , and gradient clipping at norm .
 The learning rate peaks at after a linear warm-up covering the first five percent of steps and then follows a linear decay.
@@
 | Length # Feedback | 34.6 |
-4.2 Multi-faceted Evaluation
-Human Evaluation.
-After fine-tuning the language models with the assignment–feedback pairs, to test the performance of SEFL, we do human evaluation.
-We randomly sample 150 instances from the validation set.
+4.2. Multi-faceted Evaluation
+4.2.1. Human Evaluation
+After fine-tuning the language models with the assignment-feedback pairs and testing the performance of SEFL, we conduct a human evaluation on a total of 450 instances with three annotators, following a similar approach to SoctraticLM Liu et al.
+(2024b), which annotated around 1,000 dialogues with 10 annotators.
+We randomly sample the subset of instances from the validation set.
 For each item, both the original instruction tuned model (A) and the model further fine-tuned with SEFL (B) produce feedback.
@@
 For each item they read the original prompt, the student submission, and the two candidate feedback drafts from the non-tuned and SEFL-tuned model.
-Then, they select the feedback, from model A or B, that is better based on four base criteria:
-Accuracy. The generated feedback text refers to concrete strengths and weaknesses in the student work and avoid superficial remarks.
-Actionability. Suggestions are clear, specific, and realistic for a student to apply.
-Conciseness. Wording is brief and focused, with little repetition.
-Tone. Language stays constructive and professional while recognizing good elements.
+Then, they select the feedback from model A or B that is better based on four base criteria:
+Accuracy: The generated feedback text focuses on concrete strengths and weaknesses in the student’s work, avoiding superficial remarks.
+Actionability: Suggestions are clear, specific, and realistic for a student to apply.
+Conciseness: Wording is brief and focused, with little repetition.
+Tone: Language stays constructive and professional while recognizing good elements.
 Raters were reminded to value efficiency over length, to prefer targeted advice over general principles, and to ignore formatting tricks.
 They recorded their choice as A or B and could leave an optional free-text comment.
-With this we calculate win rate (i.e., the percentage of choosing one feedback text over the other).
-This has become a de facto standard to evaluate long form text against each other (e.g., Rafailov et al.
-2023).
+With this, we calculate the win rate (i.e., the percentage of choosing one feedback text over the other).
+This has become a de facto standard to evaluate long-form text against each other (e.g., Rafailov et al., 2023).
 Each row took at most ten minutes, and the guidelines stressed taking regular breaks to sustain attention.
-We deliberately remove the tie option because a forced choice gives more informative labels and reduces hesitation, while a separate checkbox still lets raters mark assignmentanswerfeedback tuples as unrelated or nonsensical.
+We deliberately remove the tie option because a forced choice provides more informative labels and reduces hesitation.
+At the same time, a separate checkbox still allows raters to mark assignment answer feedback tuples as unrelated or nonsensical.
 This happened around 12% of the time, especially in the smaller, less capable models.
-The full annotation guidelines reported in the supplementary material in the code repository.
-Our human raters are in the aged 20–40 and from Europe.
-One identifies as female and the other two identifies as male.
-One female and male have a background in Computer Science and one male in Education Engineering.
-All have extensive experience in teaching and supervision or being taught and supervised, they all work in higher education (at different levels, i.e., research assistant and assistant professors) with near-native English proficiency.
-LLM-as-a-Judge.
-We also evaluate the fine-tuned models’ output using a LLM-as-a-judge framework, a method gaining traction as a method for evaluating free text output (Liu et al.
-2023; Zheng et al.
-2024; Chen et al.
-2023; Verga et al.
-2024; Törnberg 2023; Naismith, Mulcaire, and Burstein 2023; Gilardi, Alizadeh, and Kubli 2023; Kocmi and Federmann 2023; Huang et al.
-2024; Gu et al.
-2024; Falk et al.
-2025).
-The same 150 random instances are rated by the four LLMs, namely GPT-4o (Hurst et al.
-2024), Claude3.5-Sonnet, Command-R+, and DeepSeek-V3 (Liu et al.
-2024a).
-For the closed-source models’ LLM-as-a-judge experiments, we use their respective APIs and the total costs were approximately 10 USD.
-For every example the judge model receives the assignment prompt together with the two candidate feedback drafts.
-It is asked to decide which draft is better on the same four base criteria as the humans.
+The full annotation guidelines are reported in the supplementary material (Table˜7; Appendix˜B).
+Annotator Demographics.
+Our raters are aged 20-40 and from Europe.
+One identifies as female and the other two identify as male.
+One female and one male have a background in Computer Science, and one male has a background in Engineering Education.
+All have extensive experience in teaching and supervision, or being taught and supervised; they all work in higher education (at various levels, such as research assistant and assistant professors) with near-native English proficiency.
+4.2.2. LLM-as-a-Judge
+We also evaluate the fine-tuned models’ output using a LLM-as-a-judge framework, a method gaining traction for evaluating free text output Liu et al.
+(2023); Zheng et al.
+(2024); Chen et al.
+(2023); Verga et al.
+(2024); Törnberg (2023); Naismith et al.
+(2023); Gilardi et al.
+(2023); Kocmi and Federmann (2023); Huang et al.
+(2024); Gu et al.
+(2024); Falk et al.
+(2025).
+The three LLMs rate the same 450 random instances, namely GPT-4o Hurst et al.
+(2024), Claude3.5-Sonnet, and DeepSeek-V3 Liu et al.
+(2024a).
+For the LLM-as-a-judge experiments using closed-source models, we used the respective APIs, and the total cost was approximately 10 USD.
+For every example, the judge model receives the assignment prompt together with the two candidate feedback drafts.
+It is asked to decide which draft is better on the same four base criteria as the human annotators.
 The instruction forbids numeric grades or explanations and requires the judge to output exactly one character, A or B, producing a clean pairwise preference label.
-We report the full prompt in the supplementary material in the code repository.
-5 Results
-In Figure 2, we show the win rates of models fine-tuned with SEFL vs.
+5. Results
+In Figure˜2, we show the win rates of models fine-tuned with SEFL vs.
 their non-tuned version, evaluated by both humans and LLM-based judges.
 A value above 50% indicates that the SEFL-tuned models are preferred over their original versions.
-We show an example of the feedback in Figure 4, where we depict the abridged prompt and feedback by two tuned an non-tuned models.
+We show an example of the feedback in Figure˜4, where we depict the abridged prompt and feedback by a tuned and non-tuned model.
 Human Assessment.
-Overall, human rater evaluations in Figure 2 (H#) show that the SEFL-tuned models often attain high win rates, surpassing 90% in several cases with respect to the smaller models.
+Overall, human rater evaluations in Figure˜2 (H#) show that the SEFL-tuned models often achieve high win rates, surpassing 90% in several cases compared to the smaller models.
 The human annotators differed in their views on the 8B model’s output quality; however, they generally converged on the observation that the fine-tuned 14B model produces superior feedback compared to its original version.
@@
 In over 75% of cases, they confirmed the alignment between assignment, student response, and the feedback given, showing positive contextual relevance.
-LLM-as-a-Judge Results.
-For the LLM-as-a-judge evaluations (J#) in the same figure, we observe notable differences in win rates depending on the model and scale.
+LLM-as-a-Judge Evaluation.
+For the LLM-as-a-judge evaluations (J#) in the same figure, we observe several differences in win rates depending on the model and scale.
 The results largely mirror the human assessment trend up to the 3B scale.
-The results from the four LLM judges (J1: GPT-4o, J2: Claude-3.5-Sonnet, J3: Command-R+, J4: Deepseek-v3) reveal that SEFL-tuned models show varying levels of performance relative to their vanilla counterparts.333Models are picked based on their recency and performance on RewardBench (Lambert et al.
-2024), JudgeBench (Tan et al.
-2024), and JudgeArena (AtlaAI 2025).
-For instance, Qwen2.5-0.5B achieved the highest win rates across all four judges (62% on J3), indicating a consistent preference for the fine-tuned version.
-In contrast, larger models such as Llama-3.1-8B and Qwen2.5-14B exhibit lower win rates, particularly on J3 (16% and 10%, respectively), suggesting that fine-tuning with SEFL may yield diminishing returns or challenges at larger scales.
-The full judge prompt can be found in the supplementary material.
-Agreement.
-In Figure 3, we present the pairwise Cohen’s values (Cohen 1960) computed between each language-model judge and the human raters, to observe whether humans and LLM judges agree on which model gives better feedback.
-The agreement among humans was moderate to substantial: H1 and H3 reached , H1 and H2 , and H2 and H3 (Landis and Koch 1977).
+The results from the three LLM judges (J1: GPT-4o, J2: Claude-3.5-Sonnet, J3: Deepseek-v3) reveal that SEFL-tuned models show varying levels of performance relative to their vanilla counterparts.444Models are picked based on their performance on RewardBench Lambert et al.
+(2024), JudgeBench Tan et al.
+(2024), and JudgeArena AtlaAI (2025).
+For instance, Qwen2.5-0.5B achieved the highest win rates across all three judges, indicating a consistent preference for the fine-tuned version.
+As the parameter size increases, the difference in win rate decreases (e.g., 58% on average for 8B).
+The full judge prompt will be released in the supplementary material.
+Human and Model Agreements.
+In Figure˜3, we present the pairwise Cohen’s values Cohen (1960) computed between each LLM judge and the human rater, to observe whether humans and LLM judges agree on which model gives better feedback.
+The agreement among humans was moderate to substantial: H1 and H3 reached , H1 and H2 , and H2 and H3 Landis and Koch (1977).
 Among the models, Claude aligns most closely with both the other judges and the humans; Deepseek follows, and GPT-4o shows the weakest match.
-Across all model and human pairs the coefficients range from to , which underlines the subjectivity of feedback evaluation.
-Command R+ sits at the lower extreme, returning values between and when compared with human raters and with other judges, which signals virtually no agreement.
-This signals that the human experts mainly agree on the quality of feedback, but for LLM judges this can differ.
-Overall, there are improvements to be made in terms of agreement between LLMs and humans in terms of feedback quality.
-[…]
+Across all model and human pairs, the numbers range from 0.17 to 0.58, which shows the subjectivity of feedback evaluation.
+This suggests that human experts generally agree on the quality of feedback, but for LLM judges, this may vary.
+Overall, we see an opportunity for improvements in terms of agreement between LLMs and humans regarding feedback quality.
+| Model | H1 sample | H2 sample | H3 sample |
 | 0.5B | “The feedback repeats the assignment rather than pointing out issues.” | “A is concise but misses important detail.” | “Comments do not reflect missing parts of the answer.” |
@@
 | 14B | “Feedback is thorough but wording is opaque.” | “Long sentences reduce clarity.” | “Some comments repeat earlier points.” |
-| 14B-SEFL | “Clear and student friendly while retaining depth.” | “Uses short sentences and connects advice to rubric.” | “Groups comments by theme which avoids repetition.” |
+| 14B-SEFL | “Clear and student-friendly while retaining depth.” | “Uses short sentences and connects advice to the rubric.” | “Groups comments by theme, which avoids repetition.” |
 | Aspect | Base Model | SEFL-tuned Model | Preferred |
@@
 | Actionability | “Model A also provides partial solution.” | “Model B is not accurate and provides an answer instead of feedback.” | Base |
-6 Discussion
-6.1 Human Qualitative Insights
-In addition to the win rates in Figure 2, our human annotators provided rich qualitative feedback on the model outputs, which we show in Table 3.
-Generally, on the critical side, they noted that if a student answer is too short or incomplete, neither model explicitly flags the missing details.
+6. Discussion
+6.1. Human Qualitative Insights
+In addition to the win rates in Figure˜2, our human annotators provided rich qualitative feedback on the model outputs, which we show in Table˜4.
+Generally, on the critical side, they noted that if a student’s answer is too short or incomplete, neither model explicitly flags the missing details.
 More specifically, Qwen2.5-0.5B was praised for clarity and concision, whereas Llama-3.2-3B tended to repeat assignment details without offering actionable guidance.
@@
 Overall, although Qwen2.5-14B achieved high win rates (94, 77, 81 across three annotators), these insights suggest that even top-performing models could improve in error detection, tone refinement, and contextual sensitivity.
-To further quantify, in Figure 5, we plot for each model the net balance of optional rater comments in five qualitative categories: Actionability (AC), Goal Orientation (GO), User Friendliness (UF), Consistency (CO) and Student Autonomy (AY) (Carless et al.
-2011; Wiggins 2012).
+To further quantify, in Figure˜5, we plot for each model the net balance of optional rater comments in five qualitative categories: Actionability (AC), Goal Orientation (GO), User Friendliness (UF), Consistency (CO), and Student Autonomy (AY) Carless et al.
+(2011); Wiggins (2012).
 Squares denote the base models and circles the SEFL-tuned variants.
-Horizontal whiskers give bootstrap Wilson 95% confidence intervals for the net balance.
+Horizontal whiskers provide 95% confidence intervals for the net balance using the bootstrap method.
 We compute these intervals on the proportion of positive remarks and then transform them to the net scale via , where is the total number of comments and the positive proportion.
 Annotators added comments only when a response stood out, so the plot reveals both the direction and the strength of impressions.
-We show several examples of annotations in Table 4.
-The annotators also had a choice if the feedback was non-nonsensical, which happened around 12% of the time mainly with the smaller sized models.
+We show several examples of annotations in Table˜5.
+The annotators also had the option to reject nonsensical feedback, which occurred around 12% of the time, primarily with the smaller-sized models.
 The two smallest models, Qwen2.5-0.5B and Llama-3.2-1B, receive more negative than positive remarks on Consistency and Goal Orientation, matching earlier findings that they sometimes drift from the student answer or overlook core requirements.
 Llama-3.1-8B shows the only clearly positive balance in Actionability and User Friendliness, yet its interval for Consistency still lies below zero.
-Qwen2.5-14B gains more favorable notes on tone and clarity than the smaller models but still shows a negative alignment gap.
-With respect to Student Autonomy, there are mostly neutral comments.
-| Method/Dataset | J1 | J2 | J3 | J4 |
-| Book2Dial | 43 | 37 | 60 | 27 |
-| SEFL | 57 | 63 | 40 | 73 |
-6.2 LLM-as-a-Judge
-We used LLM judges to rate the feedback generated by SEFL-tuned models against their vanilla counterparts.
-This provides a rapid, scalable way to measure feedback quality, reducing the need for extensive human annotation.
+Qwen2.5-14B gains more favorable notes on tone and clarity than the smaller models, but still shows a negative alignment gap.
+Regarding Student Autonomy, there are mostly neutral comments.
+We show an example of the full output feedback of Qwen2.5-0.5B in Figure˜8 (Appendix˜C).
+| Method/Dataset | J1 | J2 | J3 |
+| Book2Dial | 43 | 37 | 27 |
+| SEFL | 57 | 63 | 73 |
+6.2. LLM-as-a-Judge
+We used LLM judges to rate the feedback generated by SEFL-tuned models against their untuned counterparts.
+This provides a scalable method for measuring feedback quality, thereby reducing the need for extensive human annotation.
 We let the LLM judges rate the same examples as the humans annotated.
-As shown in Figure 2, three out of four LLM judges consistently favored SEFL-tuned Qwen2.5-0.5B, Llama-3.2-1B, and Llama-3.2-3B.
-With Command-R, we notice that it performs worse than GPT-4o and Claude-3.5-Sonnet on JudgeArena.
-This indicates that the misalignment might have to do with instruction following capabilities of these models, this is highlighted by Command-R being weaker in Chatbot Arena (Zheng et al.
-2024).
-Nonetheless, we see it as a practical first step for large-scale feedback comparisons in educational contexts.
-We recommend supplementing LLM-based assessments with targeted human evaluations for more granular insights, possibly aligning more with instructional objectives or possibly even fine-tuning LLMs with rubrics for better judges of long form text (Kim et al.
-2024).
-6.3 Comparison to Prior Work
-The work that is closest to ours is Book2Dial (Wang et al.
-2024a).
+As shown in Figure˜2, 3 out of 3 LLM judges consistently favored SEFL-tuned Qwen2.5-0.5B, Llama-3.2-1B, and Llama-3.2-3B.
+We see it as a practical first step for large-scale feedback comparisons in educational contexts.
+We recommend supplementing LLM-based assessments with targeted human evaluations for more granular insights, possibly aligning more with instructional objectives or even fine-tuning LLMs with rubrics for better judgment of long-form text Kim et al.
+(2024).
+6.3. Comparison to Prior Work
+The work that is most closely related to ours is Book2Dial Wang et al.
+(2024a).
 It is a framework that turns textbooks into synthetic conversations between a student model and a teacher model.
-The student only sees high-level cues such as section titles or key concepts, while the teacher has full access to the source passage, prompting a question-answer exchange that stays aligned with the book content.
-Instead, we focus on any type of assignment not limited to textbooks.
+The student only sees high-level cues, such as section titles or key concepts, while the teacher has full access to the source passage, prompting a question-and-answer exchange that remains aligned with the book content.
+Instead, we focus on any assignment, not limited to textbooks.
 To compare the two methods, we further fine-tune Qwen2.5-14B with the Book2Dial data in the same way as our SEFL-tuned version.
-We use the existing Book2Dial data consisting of 889 dialogues, we preprocess the data such that there is only a one-turn conversation, resulting in 5.3K conversation pairs.
-We then run the fine-tuned model over the same samples evaluated in Subsection 4.2 and evaluate them with the same 4 judges and evaluation criteria.444At the time of writing, we ran out of funding for the human annotators and thus compare here only with LLM-as-a-Judge.
-Table 5 shows that three out of four judges prefer SEFL, yielding an average win rate of 58%.
-Only command-r-plus leans the other way with a 60% win rate for Book2Dial, but an average of 42% win rate.
-The finding is consistent with its earlier drift of seemingly failing to instruction follow.
-The results confirm that SEFL produces higher quality feedback than the textbook based dialogues in Book2Dial.
-7 Conclusion
+We utilize the existing Book2Dial data, which consists of 889 dialogues, and preprocess the data to ensure that each conversation is one-turn, resulting in 5,300 conversation pairs.
+We then run the fine-tuned model over the same samples evaluated in Subsection 4.2 and assess them with the same 3 judges and evaluation criteria.555At the time of writing, we ran out of funding for the human annotators and thus compare here only with LLM-as-a-Judge.
+Table˜6 shows that 3 out of 3 judges prefer SEFL, yielding an average win rate of 64%.
+The results confirm that SEFL produces higher quality feedback than the textbook-based dialogues in Book2Dial.
+7. Conclusion
 We introduced SEFL, a framework that simulates teacherstudent interactions via two-agent LLMs to generate synthetic data for fine-tuning smaller models.
-Empirical and qualitative results show the potential of this framework for enhancing feedback in educational settings.
+In this work, we aimed to explore how synthetic teacher-student interactions generated by LLMs can be leveraged for scalable feedback.
+We address this by demonstrating that these interactions, when generated through SEFL, can be used as high-quality fine-tuning data to enhance the feedback capabilities of smaller, more efficient language models.
 This approach yields concise, context-sensitive feedback that often surpasses original instruction-tuned models under both LLM-as-a-judge and human evaluations.
-Yet human insights remain indispensable for capturing nuances like clarity and tone.
+Yet, human insights remain indispensable for capturing nuances such as clarity and tone.
 SEFL provides a promising avenue for immediate, personalized feedback at scale, extending beyond the educational domain.
-Limitations.
-We acknowledge that SEFL relies on synthetically generated assignments and errors, and are not real student submissions, which could have implications.
+Limitations
+We acknowledge that SEFL relies on short-form, synthetically generated assignments and errors, which are not real student submissions, and this could have implications.
 Although this approach helps create large datasets, it risks producing feedback unaligned with authentic classroom contexts.
-Our evaluation also uses LLM-based judges, introducing potential biases related to each judge’s training data and objectives.
-Lastly, while we focused on short-answer tasks, longer or more domain-specific assignments may require specialized or more diverse synthetic data.
+Our evaluation also utilizes LLM-based judges, which introduces potential biases related to each judge’s training data and objectives.
+Lastly, while we focused on short-answer tasks, longer or more domain-specific assignments may require specialized or more diverse synthetic data, which we leave to future work.
 Ethical Statement
 The use of synthetic data provides an opportunity to train automated feedback systems without the constraints of privacy and consent that come from repurposing actual student assignments and teacher feedback as training data.
-However, it also raises questions about transparency and potential misuse (Lindsay et al.
-2024).
+However, it also raises questions about transparency and potential misuse Lindsay et al.
+(2024).
 For instance, malicious actors could manipulate synthetic data to disseminate misleading or biased feedback, undermining trust in educational tools.
 Users may also mistake synthetic feedback for real, expert guidance.
-Moreover, automated feedback systems risk reinforcing biases if the underlying models carry skewed training data.
+Moreover, automated feedback systems risk reinforcing biases if the underlying models are trained on skewed data.
 We believe educators and institutions should remain aware of these risks and incorporate human oversight to ensure that such systems complement, rather than replace, genuine pedagogical engagement with real teachers.
+Acknowledgements
+MZ, APD, JB, and EDL were supported by the research grant (VIL57392) from VILLUM FONDEN.
+We would like to thank the AAU-NLP group for helpful discussions and feedback on an earlier version of this article.
+MZ also received funding from the Danish Government to Danish Foundation Models (4378-00001B).
+We acknowledge the Danish e-Infrastructure Cooperation for awarding this project access (No.
+465001263; DeiC-AAU-N5-2024078 - H2-2024-18) to the LUMI supercomputer, owned by the EuroHPC Joint Undertaking, hosted by CSC (Finland) and the LUMI consortium through DeiC, Denmark.
+Bibliographical References
+- Alrajhi et al.
+(2021) Laila Alrajhi, Ahmed Alamri, Filipe Dwan Pereira, and Alexandra I Cristea.
+2021.
+Urgency analysis of learners’ comments: An automated intervention priority model for mooc.
+In Intelligent Tutoring Systems: 17th International Conference, ITS 2021, Virtual Event, June 7–11, 2021, Proceedings 17, pages 148–160.
+Springer.
+- Aslan et al.
+(2019) Sinem Aslan, Nese Alyuz, Cagri Tanriover, Sinem E Mete, Eda Okur, Sidney K D’Mello, and Asli Arslan Esme.
+2019.
+Investigating the impact of a real-time, multimodal student engagement analytics technology in authentic classrooms.
+In Proceedings of the 2019 chi conference on human factors in computing systems, pages 1–12.
+- AtlaAI (2025) AtlaAI. 2025. Judge arena. https://huggingface.co/spaces/AtlaAI/judge-arena. [Online; accessed 8-April-2025].
+- Bauer et al.
+(2023) Elisabeth Bauer, Martin Greisel, Ilia Kuznetsov, Markus Berndt, Ingo Kollar, Markus Dresel, Martin R Fischer, and Frank Fischer.
+2023.
+Using natural language processing to support peer-feedback in the age of artificial intelligence: A cross-disciplinary framework and a research agenda.
+British Journal of Educational Technology, 54(5):1222–1245.
+- Botelho et al.
+(2023) Anthony Botelho, Sami Baral, John A Erickson, Priyanka Benachamardi, and Neil T Heffernan.
+2023.
+Leveraging natural language processing to support automated assessment and feedback for student open responses in mathematics.
+Journal of computer assisted learning, 39(3):823–840.
+- Carless et al. (2011) David Carless, Diane Salter, Min Yang, and Joy Lam. 2011. Developing sustainable feedback practices. Studies in higher education, 36(4):395–407.
+- Chen et al.
+(2023) Yi Chen, Rui Wang, Haiyun Jiang, Shuming Shi, and Ruifeng Xu.
+2023.
+Exploring the use of large language models for reference-free text quality evaluation: An empirical study.
+In Findings of the Association for Computational Linguistics: IJCNLP-AACL 2023 (Findings), pages 361–374, Nusa Dua, Bali.
+Association for Computational Linguistics.
+- Cohen (1960) Jacob Cohen. 1960. A coefficient of agreement for nominal scales. Educational and psychological measurement, 20(1):37–46.
+- Conole and Oliver (2006) Grainne Conole and Martin Oliver. 2006. Contemporary perspectives in e-learning research. Routledge London.
+- Costello and Crane (2013) Jane Costello and Daph Crane. 2013. Technologies for learner-centered feedback. Open Praxis, 5(3):217–225.
+- Demszky and Hill (2023) Dorottya Demszky and Heather Hill.
+2023.
+The NCTE transcripts: A dataset of elementary math classroom transcripts.
+In Proceedings of the 18th Workshop on Innovative Use of NLP for Building Educational Applications (BEA 2023), pages 528–538, Toronto, Canada.
+Association for Computational Linguistics.
+- Falk et al.
+(2025) Jeanette Falk, Yiyi Chen, Janet Rafner, Mike Zhang, Johannes Bjerva, and Alexander Nolte.
+2025.
+How do hackathons foster creativity?
+towards ai collaborative evaluation of creativity at scale.
+In Proceedings of the 2025 CHI Conference on Human Factors in Computing Systems, CHI ’25.
+Association for Computing Machinery.
+- Ferguson (2011) Peter Ferguson. 2011. Student perceptions of quality feedback in teacher education. Assessment & evaluation in higher education, 36(1):51–62.
+- Fischer et al.
+(2020) Christian Fischer, Zachary A Pardos, Ryan Shaun Baker, Joseph Jay Williams, Padhraic Smyth, Renzhe Yu, Stefan Slater, Rachel Baker, and Mark Warschauer.
+2020.
+Mining big data in education: Affordances and challenges.
+Review of Research in Education, 44(1):130–160.
+- Gilardi et al.
+(2023) Fabrizio Gilardi, Meysam Alizadeh, and Maël Kubli.
+2023.
+Chatgpt outperforms crowd workers for text-annotation tasks.
+Proceedings of the National Academy of Sciences, 120(30):e2305016120.
+- Gu et al.
+(2024) Jiawei Gu, Xuhui Jiang, Zhichao Shi, Hexiang Tan, Xuehao Zhai, Chengjin Xu, Wei Li, Yinghan Shen, Shengjie Ma, Honghao Liu, Yuanzhuo Wang, and Jian Guo.
+2024.
+A survey on llm-as-a-judge.
+- Guerraoui et al.
+(2023) Camelia Guerraoui, Paul Reisert, Naoya Inoue, Farjana Sultana Mim, Keshav Singh, Jungmin Choi, Irfan Robbani, Shoichi Naito, Wenzhi Wang, and Kentaro Inui.
+2023.
+Teach me how to argue: A survey on NLP feedback systems in argumentation.
+In Proceedings of the 10th Workshop on Argument Mining, pages 19–34, Singapore.
+Association for Computational Linguistics.
+- Hattie (2008) John Hattie. 2008. Visible learning: A synthesis of over 800 meta-analyses relating to achievement. routledge.
+- Huang et al.
+(2024) Fan Huang, Haewoon Kwak, Kunwoo Park, and Jisun An.
+2024.
+ChatGPT rates natural language explanation quality like humans: But on which scales?
+In Proceedings of the 2024 Joint International Conference on Computational Linguistics, Language Resources and Evaluation (LREC-COLING 2024), pages 3111–3132, Torino, Italia.
+ELRA and ICCL.
+- Hurst et al.
+(2024) Aaron Hurst, Adam Lerer, Adam P Goucher, Adam Perelman, Aditya Ramesh, Aidan Clark, AJ Ostrow, Akila Welihinda, Alan Hayes, Alec Radford, et al.
+2024.
+Gpt-4o system card.
+arXiv preprint arXiv:2410.21276.
+- Ke and Ng (2019) Zixuan Ke and Vincent Ng. 2019. Automated essay scoring: A survey of the state of the art. In IJCAI, volume 19, pages 6300–6308.
+- Kim et al.
+(2022) Gangwoo Kim, Sungdong Kim, Kang Min Yoo, and Jaewoo Kang.
+2022.
+Generating information-seeking conversations from unlabeled documents.
+In Proceedings of the 2022 Conference on Empirical Methods in Natural Language Processing, pages 2362–2378, Abu Dhabi, United Arab Emirates.
+Association for Computational Linguistics.
+- Kim et al.
+(2024) Seungone Kim, Juyoung Suk, Shayne Longpre, Bill Yuchen Lin, Jamin Shin, Sean Welleck, Graham Neubig, Moontae Lee, Kyungjae Lee, and Minjoon Seo.
+2024.
+Prometheus 2: An open source language model specialized in evaluating other language models.
+arXiv preprint arXiv:2405.01535.
+- Kocmi and Federmann (2023) Tom Kocmi and Christian Federmann.
+2023.
+Large language models are state-of-the-art evaluators of translation quality.
+In Proceedings of the 24th Annual Conference of the European Association for Machine Translation, pages 193–203, Tampere, Finland.
+European Association for Machine Translation.
+- Kwon et al. (2024) Soonwoo Kwon, Sojung Kim, Minju Park, Seunghyun Lee, and Kyuseok Kim. 2024. Biped: Pedagogically informed tutoring system for esl education. arXiv preprint arXiv:2406.03486.
+- Lambert et al.
+(2024) Nathan Lambert, Valentina Pyatkin, Jacob Morrison, LJ Miranda, Bill Yuchen Lin, Khyathi Chandu, Nouha Dziri, Sachin Kumar, Tom Zick, Yejin Choi, et al.
+2024.
+Rewardbench: Evaluating reward models for language modeling.
+arXiv preprint arXiv:2403.13787.
+- Landis and Koch (1977) J Richard Landis and Gary G Koch. 1977. The measurement of observer agreement for categorical data. biometrics, pages 159–174.
+- Li et al.
+(2023) Guohao Li, Hasan Hammoud, Hani Itani, Dmitrii Khizbullin, and Bernard Ghanem.
+2023.
+Camel: Communicative agents for" mind" exploration of large language model society.
+Advances in Neural Information Processing Systems, 36:51991–52008.
+- Liang et al.
+(2024) Weixin Liang, Yuhui Zhang, Hancheng Cao, Binglu Wang, Daisy Yi Ding, Xinyu Yang, Kailas Vodrahalli, Siyu He, Daniel Scott Smith, Yian Yin, et al.
+2024.
+Can large language models provide useful feedback on research papers?
+a large-scale empirical analysis.
+NEJM AI, 1(8):AIoa2400196.
+- Lindsay et al. (2024) Euan D Lindsay, Mike Zhang, Aditya Johri, and Johannes Bjerva. 2024. The responsible development of automated student feedback with generative ai.
+- Liu et al.
+(2024a) Aixin Liu, Bei Feng, Bing Xue, Bingxuan Wang, Bochao Wu, Chengda Lu, Chenggang Zhao, Chengqi Deng, Chenyu Zhang, Chong Ruan, et al.
+2024a.
+Deepseek-v3 technical report.
+arXiv preprint arXiv:2412.19437.
+- Liu et al.
+(2024b) Jiayu Liu, Zhenya Huang, Tong Xiao, Jing Sha, Jinze Wu, Qi Liu, Shijin Wang, and Enhong Chen.
+2024b.
+Socraticlm: exploring socratic personalized teaching with large language models.
+Advances in Neural Information Processing Systems, 37:85693–85721.
+- Liu et al.
+(2023) Yang Liu, Dan Iter, Yichong Xu, Shuohang Wang, Ruochen Xu, and Chenguang Zhu.
+2023.
+G-eval: NLG evaluation using gpt-4 with better human alignment.
+In Proceedings of the 2023 Conference on Empirical Methods in Natural Language Processing, pages 2511–2522, Singapore.
+Association for Computational Linguistics.
+- Liu et al.
+(2024c) Zhengyuan Liu, Stella Xin Yin, Geyu Lin, and Nancy F Chen.
+2024c.
+Personality-aware student simulation for conversational intelligent tutoring systems.
+arXiv preprint arXiv:2404.06762.
+- Loshchilov and Hutter (2019) Ilya Loshchilov and Frank Hutter.
+2019.
+Decoupled weight decay regularization.
+In 7th International Conference on Learning Representations, ICLR 2019, New Orleans, LA, USA, May 6-9, 2019.
+OpenReview.net.
+- Lozhkov et al. (2024) Anton Lozhkov, Loubna Ben Allal, Leandro von Werra, and Thomas Wolf. 2024. Fineweb-edu.
+- Nair et al.
+(2024) Inderjeet Jayakumar Nair, Jiaye Tan, Xiaotian Su, Anne Gere, Xu Wang, and Lu Wang.
+2024.
+Closing the loop: Learning to generate writing feedback via language model simulated student revisions.
+In Proceedings of the 2024 Conference on Empirical Methods in Natural Language Processing, pages 16636–16657, Miami, Florida, USA.
+Association for Computational Linguistics.
+- Naismith et al.
+(2023) Ben Naismith, Phoebe Mulcaire, and Jill Burstein.
+2023.
+Automated evaluation of written discourse coherence using GPT-4.
+In Proceedings of the 18th Workshop on Innovative Use of NLP for Building Educational Applications (BEA 2023), pages 394–403, Toronto, Canada.
+Association for Computational Linguistics.
+- Nicol (2007) David Nicol. 2007. E-assessment by design: using multiple-choice tests to good effect. Journal of Further and higher Education, 31(1):53–64.
+- Rafailov et al.
+(2023) Rafael Rafailov, Archit Sharma, Eric Mitchell, Christopher D Manning, Stefano Ermon, and Chelsea Finn.
+2023.
+Direct preference optimization: Your language model is secretly a reward model.
+Advances in neural information processing systems, 36:53728–53741.
+- Ramesh and Sanampudi (2022) Dadi Ramesh and Suresh Kumar Sanampudi. 2022. An automated essay scoring systems: a systematic literature review. Artificial Intelligence Review, 55(3):2495–2527.
+- Rooein and Hovy (2024) Donya Rooein and Dirk Hovy. 2024. Conversations as a source for teaching scientific concepts at different education levels. arXiv preprint arXiv:2404.10475.
+- Ross and Andreas (2024) Alexis Ross and Jacob Andreas.
+2024.
+Toward in-context teaching: Adapting examples to students’ misconceptions.
+In Proceedings of the 62nd Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), pages 13283–13310, Bangkok, Thailand.
+Association for Computational Linguistics.
+- Saito et al. (2023) Keita Saito, Akifumi Wachi, Koki Wataoka, and Youhei Akimoto. 2023. Verbosity bias in preference labeling by large language models. arXiv preprint arXiv:2310.10076.
+- Schwarz et al.
+(2018) Baruch B Schwarz, Naomi Prusak, Osama Swidan, Adva Livny, Kobi Gal, and Avi Segal.
+2018.
+Orchestrating the emergence of conceptual learning: A case study in a geometry class.
+International Journal of Computer-Supported Collaborative Learning, 13:189–211.
+- Sonkar et al. (2024) Shashank Sonkar, Kangqi Ni, Sapana Chaudhary, and Richard G Baraniuk. 2024. Pedagogical alignment of large language models. arXiv preprint arXiv:2402.05000.
+- Stahl et al.
+(2024) Maja Stahl, Leon Biermann, Andreas Nehring, and Henning Wachsmuth.
+2024.
+Exploring LLM prompting strategies for joint essay scoring and feedback generation.
+In Proceedings of the 19th Workshop on Innovative Use of NLP for Building Educational Applications (BEA 2024), pages 283–298, Mexico City, Mexico.
+Association for Computational Linguistics.
+- Suresh et al.
+(2022) Abhijit Suresh, Jennifer Jacobs, Charis Harty, Margaret Perkoff, James H.
+Martin, and Tamara Sumner.
+2022.
+The TalkMoves dataset: K-12 mathematics lesson transcripts annotated for teacher and student discursive moves.
+In Proceedings of the Thirteenth Language Resources and Evaluation Conference, pages 4654–4662, Marseille, France.
+European Language Resources Association.
+- Tan et al.
+(2024) Sijun Tan, Siyuan Zhuang, Kyle Montgomery, William Y Tang, Alejandro Cuadron, Chenguang Wang, Raluca Ada Popa, and Ion Stoica.
+2024.
+Judgebench: A benchmark for evaluating llm-based judges.
+arXiv preprint arXiv:2410.12784.
+- Törnberg (2023) Petter Törnberg. 2023. Chatgpt-4 outperforms experts and crowd workers in annotating political twitter messages with zero-shot learning. arXiv preprint arXiv:2304.06588.
+- Verga et al.
+(2024) Pat Verga, Sebastian Hofstatter, Sophia Althammer, Yixuan Su, Aleksandra Piktus, Arkady Arkhangorodsky, Minjie Xu, Naomi White, and Patrick Lewis.
+2024.
+Replacing judges with juries: Evaluating llm generations with a panel of diverse models.
+arXiv preprint arXiv:2404.18796.
+- Wang et al.
+(2024a) Junling Wang, Jakub Macina, Nico Daheim, Sankalan Pal Chowdhury, and Mrinmaya Sachan.
+2024a.
+Book2Dial: Generating teacher student interactions from textbooks for cost-effective development of educational chatbots.
+In Findings of the Association for Computational Linguistics: ACL 2024, pages 9707–9731, Bangkok, Thailand.
+Association for Computational Linguistics.
+- Wang and Demszky (2024) Rose Wang and Dorottya Demszky.
+2024.
+Edu-ConvoKit: An open-source library for education conversation data.
+In Proceedings of the 2024 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies (Volume 3: System Demonstrations), pages 61–69, Mexico City, Mexico.
+Association for Computational Linguistics.
+- Wang et al.
+(2024b) Rose E Wang, Ana T Ribeiro, Carly D Robinson, Susanna Loeb, and Dora Demszky.
+2024b.
+Tutor copilot: A human-ai approach for scaling real-time expertise.
+arXiv preprint arXiv:2410.03017.
+- Wang et al.
+(2024c) Shen Wang, Tianlong Xu, Hang Li, Chaoli Zhang, Joleen Liang, Jiliang Tang, Philip S Yu, and Qingsong Wen.
+2024c.
+Large language models for education: A survey and outlook.
+arXiv preprint arXiv:2403.18105.
+- Wiggins (2012) Grant Wiggins. 2012. Seven keys to effective feedback. Feedback, 70(1):10–16.
+- Wu et al.
+(2023) Qingyun Wu, Gagan Bansal, Jieyu Zhang, Yiran Wu, Shaokun Zhang, Erkang Zhu, Beibin Li, Li Jiang, Xiaoyun Zhang, and Chi Wang.
+2023.
+Autogen: Enabling next-gen llm applications via multi-agent conversation framework.
+arXiv preprint arXiv:2308.08155.
+- Yun et al.
+(2024) Joy Yun, Yann Hicke, Mariah Olson, and Dorottya Demszky.
+2024.
+Enhancing tutoring effectiveness through automated feedback: Preliminary findings from a pilot randomized controlled trial on sat tutoring.
+In Proceedings of the Eleventh ACM Conference on Learning@ Scale, pages 422–426.
+- Zhang et al.
+(2025) Mike Zhang, Euan Lindsay, Maj-Britt Quitzau, and Johannes Bjerva.
+2025.
+Scaling course evaluations with large language models: Semester-level digestible student feedback for program leaders.
+In Proceedings of the 53rd Annual Conference of the European Society for Engineering Education (SEFI 2025).
+- Zhang et al.
+(2024) Mike Zhang, Euan Lindsay, Frederik Bode Thorbensen, Danny Bøgsted Poulsen, and Johannes Bjerva.
+2024.
+Leveraging large language models for actionable course evaluation student feedback to lecturers.
+In Proceedings of the 52nd Annual Conference of the European Society for Engineering Education (SEFI), pages 1089–1098.
+- Zhang et al.
+(2020) Tianyi Zhang, Varsha Kishore, Felix Wu, Kilian Q.
+Weinberger, and Yoav Artzi.
+2020.
+Bertscore: Evaluating text generation with BERT.
+In 8th International Conference on Learning Representations, ICLR 2020, Addis Ababa, Ethiopia, April 26-30, 2020.
+OpenReview.net.
+- Zheng et al.
+(2024) Lianmin Zheng, Wei-Lin Chiang, Ying Sheng, Siyuan Zhuang, Zhanghao Wu, Yonghao Zhuang, Zi Lin, Zhuohan Li, Dacheng Li, Eric Xing, et al.
+2024.
+Judging llm-as-a-judge with mt-bench and chatbot arena.
+Advances in Neural Information Processing Systems, 36.
+Appendix
+Appendix A Prompts
+In Figure˜6, we show the prompts that we give to the agent models. Additionally, in Figure˜7, we show the LLM-as-a-judge that we give to the judge models.
+Appendix B Human Evaluation Guidelines
+In Table˜7, we show the annotation guidelines for the human raters to rate the model feedback. The annotators were also instructed that the data will be made publicly available.
+| Section | Details |
+| Overview |
+Your task is to evaluate pairs of feedback responses (Model A and Model B) given to student assignments. You will select which model provides better feedback according to specific criteria.
+Key Principles: • Focus on efficiency and specificity.
+• Value concise, meaningful feedback over lengthy explanations.
+• Prioritize direct, actionable suggestions.
+• Consider both content and delivery.
+Remember to take breaks; I suggest spending a maximum of 10 minutes per row.
+|
+| Sheet Information | In the table, pick the one you got assigned.
+You will see 7 columns and need to fill in columns C and F: • Appendix_assignment: What the large language model saw when generating an assignment with a possible answer.
+• Assignment: What the model generated as an assignment and answered.
+• Model A: Feedback generated by Model A.
+• Model B: Feedback generated by Model B.
+• Which is better?
+The most important part is to evaluate both feedback responses and determine which one is better, based on the assignment and answer.
+• Comments: Leave comments if needed.
+|
+| Evaluation Criteria |
+Accuracy: Does the feedback address specific strengths and weaknesses? Are comments relevant to the student work? Is the critique substantive rather than superficial?
+Actionability: Are suggestions clear and specific?
+Can students easily understand what to improve?
+Are recommendations implementable?
+Conciseness: Is the feedback brief while remaining meaningful?
+Does it avoid unnecessary elaboration?
+Is there minimal redundancy?
+Tone: Is the feedback constructive while being efficient?
+Does it balance recognition with criticism?
+Is the language professional?
+|
+| Format | Preferred Feedback Style: • Shows good understanding of the concept.
+• Uses specific examples from the text to support arguments.
+• Addresses the main question directly.
+Less Preferred Feedback Style: • Generalized or vague feedback.
+• Overly verbose or structured responses.
+• Focuses on theoretical completeness rather than practical advice.
+|
+| Scoring and Pitfalls | Scoring: 1.
+Read the original assignment carefully.
+2.
+Review both feedback responses.
+3.
+Evaluate against the criteria.
+4.
+Select the model that better aligns with the criteria as “A” or “B.” Pitfalls: • Avoid preferring longer feedback just because it’s lengthy.
+• Do not choose feedback that only lists general principles.
+• Avoid letting formatting alone affect your choice.
+|
+Appendix C Feedback Example
+In Figure˜8, we show an example of the feedback of Qwen2.5-0.5B tuned on SEFL and out-of-the-box.